### 26 Mar 2026 v1

# Optimal High-Probability Regret for Online Convex Optimization with Two-Point Bandit Feedback

Abstract
We consider the problem of Online Convex Optimization (OCO) with two-point bandit feedback in an adversarial environment. In this setting, a player attempts to minimize a sequence of adversarially generated convex loss functions, while only observing the value of each function at two points. While it is well-known that two-point feedback allows for gradient estimation, achieving tight high-probability regret bounds for strongly convex functions still remained open as highlighted by Agarwal et al. (2010). The primary challenge lies in the heavy-tailed nature of bandit gradient estimators, which makes standard concentration analysis difficult. In this paper, we resolve this open challenge by providing the first high-probability regret bound of $O(d(\log T+\log(1/\delta))/\mu)$ for $\mu$-strongly convex losses. Our result is minimax optimal with respect to both the time horizon $T$ and the dimension $d$.
1 Introduction
Online Convex Optimization (OCO) can be understood as a repeated game between a player and an adversary. At each round $t=1,\dots,T$, the player chooses an action $x_{t}$ from a fixed convex set $\mathcal{K}\subset\mathbb{R}^{d}$. Simultaneously, an adversary selects a convex loss function $\ell_{t}:\mathcal{K}\to\mathbb{R}$. The player then suffers the loss $\ell_{t}(x_{t})$, and their performance is measured by the regret:
$R_{T}=\sum_{t=1}^{T}\ell_{t}(x_{t})-\min_{x\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(x).$
While full-information OCO is well-studied, the bandit feedback setting, where only the values of $\ell_{t}$ are observed, remains a central challenge due to the lack of direct gradient information.
To bridge the gap between bandit and full-information feedback where the gradient information of the loss function $\ell_{t}(\cdot)$ can be obtained, Agarwal et al. (2010) introduced the multi-point bandit setting. In their seminal work, they demonstrated that evaluating the loss function at two randomized points allows the player to construct gradient estimators that mimic the behavior of true gradients. However, as explicitly discussed in their paper, achieving tight high-probability regret bounds for strongly convex loss functions under two-point feedback remained an unresolved challenge. Specifically, while expectation bounds were established, the heavy-tailed nature of bandit estimators against an adversary has historically made high-confidence guarantees elusive.
In this work, we resolve this open problem by providing the first high-probability regret bound for two-point feedback OCO with $\mu$-strongly convex losses. Our main result establishes that with probability at least $1-\delta$, the regret satisfies:
$R_{T}\leq O\left(\frac{d(\log T+\log(1/\delta))}{\mu}\right).$
This regret is minimax optimal both in $T$ (Abernethy et al., 2009) and dimension $d$ (Duchi et al., 2015).
Distinct from the conventional approach that directly reduces the analysis to smoothed loss functions under standard Online Gradient Descent (OGD), we propose a novel analytical framework. By departing from this reduction-based paradigm, we are able to derive a bound that is more robust to the variance of the zero-order estimator.
Furthermore, our result significantly improves the dependency on the dimension $d$. We reduce the regret from the $O(d^{2})$ factor found in prior work (Agarwal et al., 2010). This improvement is achieved by extending the sophisticated geometric and probabilistic techniques introduced by Shamir (2017), which were used to prove a regret holding in expectation, to the high-probability regime. By generalizing these techniques to accommodate high-confidence concentration, we match the information-theoretic lower bounds for this online convex optimization setting.
2 Notation and Preliminaries
2.1 Notation
We first define our notation will appear in various parts of the analysis. Let $\|\cdot\|$ denote the Euclidean norm in $\mathbb{R}^{d}$, and let $\mathcal{B}=\{\bm{x}\in\mathbb{R}^{d}:\|\bm{x}\|\leq 1\}$ be the unit ball centered at the origin. We assume that the convex set $\mathcal{K}$ is compact and has a nonempty interior. If not, we can always map $\mathcal{K}$ to a lower dimensional space. More specifically, we assume that there exist $D>0$ such that
$\mathcal{K}\subseteq D\mathcal{B}.$ (1)
We assume w.l.o.g. that the set contains 0, as we can always translate $\mathcal{K}$.
Furthermore, in this paper, we will assume that loss functions $\ell_{t}(\cdot)$ are $G$-Lipschitz continuous and $\mu$-strongly convex.
Assumption 1
We assume that each loss function $\ell_{t}$ is Lipschitz continuous, i.e., there exists a constant $G\geq 0$ such that
$|\ell_{t}(\bm{x})-\ell_{t}(\bm{y})|\leq G\|\bm{x}-\bm{y}\|,\quad\forall\bm{x},\bm{y}\in\mathcal{K},\forall t.$
Assumption 2
For $\mu\geq 0$, the function $\ell$ is called $\mu$-strongly convex on the set $\mathcal{K}$ if
$\ell(\bm{x})\geq\ell(\bm{y})+\nabla\ell(\bm{y})^{\top}(\bm{x}-\bm{y})+\frac{\mu}{2}\|\bm{x}-\bm{y}\|^{2},\quad\forall x,y\in\mathcal{K}.$
2.2 Preliminaries
We first define the smoothed loss function $\hat{\ell}_{t}(\cdot)$ as follows:
$\hat{\ell}_{t}(\bm{x})=\mathbb{E}_{\bm{v}}\left[\ell_{t}(\bm{x}+\alpha\bm{v})\right],$ (2)
where $\bm{v}$ is a uniform random vector in the unit ball $\mathcal{B}$. Then, the smoothed function $\hat{\ell}_{t}(\cdot)$ can preserve some properties of $\ell_{t}(\cdot)$ just shown as follows.
Lemma 1 (Nesterov and Spokoiny (2017))
If $\ell(\bm{x})$ is $G$-Lipschitz continuous, then $\hat{\ell}(\bm{x})$ defined in Eq. (2) is also at least $G$-Lipschitz continuous. If $\ell(\bm{x})$ is $\mu$-strongly convex, then $\hat{\ell}(\bm{x})$ defined in Eq. (2) is also at least $\mu$-strongly convex.
At the same time, we can construct an unbiased estimation of $\nabla\hat{\ell}_{t}(\bm{x}_{t})$ as follows:
$\bm{g}_{t}=d\cdot\frac{\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})}{2\alpha}\cdot\bm{u}_{t},\mbox{ where }\bm{u}_{t}\sim\mathbb{S}^{d-1},$ (3)
that is, $\bm{u}_{t}$ is uniformly drawn from the sphere of a unit ball.
Lemma 2 (Lemma 10 of Shamir (2017))
The approximate gradient $\bm{g}_{t}$ defined in Eq. (3) satisfies that $\mathbb{E}[\bm{g}_{t}\mid\bm{x}_{t}]=\nabla\hat{\ell}_{t}(\bm{x}_{t})$ and $\mathbb{E}[\|\bm{g}_{t}\|^{2}\mid\bm{x}_{t}]\leq c_{0}dG^{2}$ where $c_{0}$ is an absolute constant.
Above lemma shows that the approximate gradient $\bm{g}_{t}$ is an unbiased estimation of $\nabla\hat{\ell}_{t}(\cdot)$ and has a bounded variance linearly depending on the dimension $d$. Accordingly, we can use the approximate gradient $\bm{g}_{t}$ to achieve the online gradient descent just as shown in Algorithm 1.
Lemma 3 (Lemma 2 of Agarwal et al. (2010))
For any $\bm{x}\in\mathcal{K}$, it holds that
$\sum_{t=1}^{T}\frac{\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})+\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})}{2}-\sum_{t=1}^{T}\ell_{t}(\bm{x})$ (4)
$\leq$ $\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\sum_{t=1}^{T}\hat{\ell}_{t}((1-\xi)\bm{x})+3TG\alpha+TGD\xi.$
Algorithm 1 Online Gradient Descent with two queries per round
1:Input: Learning rates $\eta_{t}$, exploration parameter $\alpha$ and shrinkage coefficient $\xi$.
2: Set $\bm{x}_{1}=0$
3:for $t=1,\dots,T$ do
4:  Pick a unit vector $\bm{u}_{t}$ uniformly at random from the sphere of a unit ball.
5:  Observe $\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})$ and $\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})$.
6:  Construct the approximate gradient $\bm{g}_{t}$ as Eq. (3).
7:  Update $\bm{x}_{t+1}=\Pi_{(1-\xi)\mathcal{K}}(\bm{x}_{t}-\eta_{t}\bm{g}_{t})$.
8:end for
3 Main Results
We first give the main result in this paper. Then, we will provide the detailed proofs.
Theorem 4
Suppose the loss function $\ell_{t}(\cdot)$ is $\mu$-strongly convex and $G$-Lipschitz continuous. The constraint set $\mathcal{K}$ is convex and satisfies Eq. (1). Let the sequence $\{\bm{x}_{t}\}$ be generated by Algorithm 1. Given $0<\delta<1$, by setting the step size $\eta_{t}=\frac{2}{\mu t}$, $\alpha=\frac{dG\log\frac{1}{\delta}}{T\mu}$ and $\xi=(d+1)\log\frac{1}{\delta}/T$, then with a probability at least $1-\delta$, it holds that for any fixed $\bm{x}\in\mathcal{K}$
$\sum_{t=1}^{T}\frac{\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})+\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})}{2}-\sum_{t=1}^{T}\ell_{t}(\bm{x})$ (5)
$\leq$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+\frac{(32c_{0}+3)dG^{2}\log\frac{1}{\delta}}{\mu}+\frac{7(d+1)GD\log\frac{1}{\delta}}{3},$
where $K$ and $c_{0}$ are two absolute constants.
Theorem 4 shows that with properly chosen parameters, Algorithm 1 can achieve a regret with a probability at least $1-\delta$
$R_{T}\leq O\left(\frac{dG^{2}(\log T+\log(1/\delta))}{\mu}\right).$ (6)
Above regret matches the one of Agarwal et al. (2010) which holds in expectation except that our regret bound has a linear dependence on the dimension while Agarwal et al. (2010) quadratically depends on the dimension. Thus, our result resolves the open problem left by Agarwal et al. (2010) whether one can achieve a $O(\log T)$ regret with high-probability for strongly convex loss functions under two-point feedback. Furthermore, the regret in Eq. (6) is the first one which achieves the minimax optimal both in $T$ (Abernethy et al., 2009) and dimension $d$ (Duchi et al., 2015).
3.1 Detailed Proofs
Our proof will revisit the proof of online convex optimization which achieves the $O(\log T)$ regret for the strongly convex function (Hazan et al., 2007). This is the key step to achieve our two-point feed-back algorithm with $O(\log T)$ regret holding in a high probability.
Lemma 5
Suppose the loss function $\ell_{t}(\cdot)$ is $\mu$-strongly convex. Let the sequence $\{\bm{x}_{t}\}$ be generated by Algorithm 1. For any $\bm{x}\in(1-\xi)\mathcal{K}$, it holds that
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq$ $\sum_{t=1}^{T}\frac{\|\bm{x}_{t}-\bm{x}\|^{2}-\|\bm{x}_{t+1}-\bm{x}\|^{2}}{2\eta_{t}}+\sum_{t=1}^{T}\frac{\eta_{t}}{2}\|\bm{g}_{t}\|^{2}$ (7)
$+\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle-\frac{\mu}{2}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}.$
Proof First, by the property of projection and the convexity of $\mathcal{K}$, we have
$\|\bm{x}_{t+1}-\bm{x}\|^{2}=$ $\left\|\Pi_{(1-\xi)\mathcal{K}}(\bm{x}_{t}-\eta_{t}\bm{g}_{t})\right\|^{2}\leq\left\|\bm{x}_{t}-\eta_{t}\bm{g}_{t}\right\|^{2}$
$=$ $\|\bm{x}_{t}-\bm{x}\|^{2}-2\eta_{t}\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle+\eta_{t}^{2}\|\bm{g}_{t}\|^{2}.$
Thus, we can obtain that
$\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle\leq\frac{\|\bm{x}_{t}-\bm{x}\|^{2}-\|\bm{x}_{t+1}-\bm{x}\|^{2}}{2\eta_{t}}+\frac{\eta_{t}}{2}\|\bm{g}_{t}\|^{2}.$ (8)
At the same time, by the $\mu$-strong convexity of $\hat{\ell}_{t}(\bm{x})$ shown in Lemma 1, we have
$\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq\langle\nabla\hat{\ell}_{t}(\bm{x}_{t}),\bm{x}_{t}-\bm{x}\rangle-\frac{\mu}{2}\|\bm{x}_{t}-\bm{x}\|^{2}.$
Thus,
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq$ $\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t}),\bm{x}_{t}-\bm{x}\rangle-\frac{\mu}{2}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}$
$=$ $\sum_{t=1}^{T}\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle+\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle-\frac{\mu}{2}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}$
[…]
$+\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle-\frac{\mu}{2}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}.$
By taking expectation to Eq. (7), we can obtain that
$\sum_{t=1}^{T}\mathbb{E}[\hat{\ell}_{t}(\bm{x}_{t})]-\hat{\ell}_{t}(\bm{x})\leq\sum_{t=1}^{T}\frac{\|\bm{x}_{t}-\bm{x}\|^{2}-\|\bm{x}_{t+1}-\bm{x}\|^{2}}{2\eta_{t}}+\sum_{t=1}^{T}\frac{\eta_{t}}{2}\mathbb{E}[\|\bm{g}_{t}\|^{2}]-\frac{\mu}{2}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}.$
Then, by setting step size $\eta_{t}=\frac{1}{\mu t}$ and upper bounding $\mathbb{E}[\|\bm{g}_{t}\|^{2}]\leq O(d^{2})$, then following the proof of Hazan et al. (2007), we can obtain the regret $O(d^{2}\log T)$ in expectation achieved in Agarwal et al. (2010).
To achieve a high-probability probability, one has to bound the extra term $\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle$ while it will disappear if we only need a result holding in expectation. However, it will incur an additional term of $O(\sqrt{T})$ by directly using the Hoffding-Azuma’s inequality (Agarwal et al., 2010).
Next, the following lemma gives a novel upper bound of $\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle$ holding with a high probability by using the Freedman’s inequality. This is a key step to overcome the additional term of $O(\sqrt{T})$ caused by directly using the Hoffding-Azuma’s inequality.
Lemma 6
Suppose the loss function $\ell_{t}(\bm{x})$ is $G$-Lipschitz continuous and the convex set $\mathcal{K}$ satisfies Eq. (1). Given $0<\delta<1$, then with a probability at least $1-\delta$, it holds that
$\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle\leq\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu},$ (9)
where $c_{0}$ is an absolute constant.
Proof Let us denote
$Z_{t}:=\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle$
By Lemma 2, we can obtain that
$\mathbb{E}[Z_{t}\mid\bm{x}_{t}]=0.$
Furthermore,
$|Z_{t}|\leq$ $\|\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t}\|\cdot\|\bm{x}_{t}-\bm{x}\|\leq(\|\nabla\hat{\ell}_{t}(\bm{x}_{t})\|+\|\bm{g}_{t}\|)\cdot\|\bm{x}_{t}-\bm{x}\|$
$\leq$ $(G+dG)\cdot 2D=2(d+1)GD,$
where the last inequality is because the loss function $\ell_{t}(\bm{x})$ is $G$-Lipschitz continuous, Lemma 1, Eq. (1), and
[…]
Next, we have
$\mathbb{E}[Z_{t}^{2}\mid\bm{x}_{t}]=$ $\mathbb{E}[\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\mid\bm{x}_{t}]$
$=$ $\mathbb{E}\left[\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\mid\bm{x}_{t}\right]+\langle\nabla\hat{\ell}_{t}(\bm{x}_{t}),\bm{x}_{t}-\bm{x}\rangle^{2}$
$-2\mathbb{E}\left[\langle\nabla\hat{\ell}_{t}(\bm{x}_{t}),\bm{x}_{t}-\bm{x}\rangle\cdot\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle\mid\bm{x}_{t}\right]$
$=$ $\mathbb{E}\left[\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\mid\bm{x}_{t}\right]-\langle\nabla\hat{\ell}_{t}(\bm{x}_{t}),\bm{x}_{t}-\bm{x}\rangle^{2}$
$\leq$ $\mathbb{E}\left[\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\mid\bm{x}_{t}\right]\leq\mathbb{E}[\|\bm{g}_{t}\|^{2}]\cdot\|\bm{x}_{t}-\bm{x}\|^{2}$
$\leq$ $c_{0}dG^{2}\|\bm{x}_{t}-\bm{x}\|^{2},$
where the third equality and the last inequality are both because of Lemma 2.
Thus, we have
$\sum_{t=1}^{T}\mathbb{E}[Z_{t}^{2}\mid\bm{x}_{t}]\leq c_{0}dG^{2}\cdot\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}.$
Then, by Lemma 11 (Freedman’s inequality) with $R=2(d+1)GD$ and $\sigma^{2}=c_{0}dG^{2}\cdot\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}$, we can obtain that with a probability at least $1-\delta$, it holds that
$\sum_{t=1}^{T}Z_{t}\leq$ $\frac{2R\log\frac{1}{\delta}}{3}+\sqrt{2\sigma^{2}\log\frac{1}{\delta}}\leq\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\sqrt{2c_{0}dG^{2}\log\frac{1}{\delta}\cdot\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}}$
$\leq$ $\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}.$
The bound in Eq. (9) is important to achieve the $O(\log T)$ regret with high probability. The bound is upper bounded by $\frac{\mu}{4}\sum_{t=1}^{T}\left\|\bm{x}_{t}-\bm{x}\right\|^{2}$. At the same time, Eq. (7) has the terms
$\sum_{t=1}^{T}\frac{\|\bm{x}_{t}-\bm{x}\|^{2}-\|\bm{x}_{t+1}-\bm{x}\|^{2}}{2\eta_{t}}-\frac{\mu}{2}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}.$
With a properly chosen step size $\eta_{t}=2/(\mu t)$ instead of $1/(\mu t)$ chosen in Agarwal et al. (2010) and Hazan et al. (2007), we can obtain the following regret where no extra $\sqrt{T}$ term shows.
Lemma 7
Suppose the loss function $\ell_{t}(\cdot)$ is $\mu$-strongly convex and $G$-Lipschitz continuous. Let the sequence $\{\bm{x}_{t}\}$ be generated by Algorithm 1. Given $0<\delta<1$, and setting the step size $\eta_{t}=2/(\mu t)$, then with a probability at least $1-\delta$, it holds that for any $\bm{x}\in(1-\xi)\mathcal{K}$:
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}.$ (10)
Proof We have
[…]
$+\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle-\frac{\mu}{2}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}$
[…]
$+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}-\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}$
$=$ $\sum_{t=1}^{T}\frac{\mu t\left(\|\bm{x}_{t}-\bm{x}\|^{2}-\|\bm{x}_{t+1}-\bm{x}\|^{2}\right)}{4}-\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}$
$+\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu},$
where the last equality is because of the step size $\eta_{t}=\frac{2}{\mu t}$.
Furthermore,
$\sum_{t=1}^{T}\frac{\mu t\left(\|\bm{x}_{t}-\bm{x}\|^{2}-\|\bm{x}_{t+1}-\bm{x}\|^{2}\right)}{4}-\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}$
$=$ $\mu\sum_{t=1}^{T}\frac{(t-1)\|\bm{x}_{t}-\bm{x}\|^{2}-t\|\bm{x}_{t+1}-\bm{x}\|^{2}}{4}$
$=$ $\mu\cdot\frac{-T\|\bm{x}_{T+1}-\bm{x}\|^{2}}{4}$
$\leq$ $0.$
Combining above results, we can obtain the final result.
Finally, we will show $\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}$ is upper bounded by $O(d\log T)$. This step is the key step that guarantees Algorithm 1 can achieve a regret linear dependence on the dimension. This is achieved by extending the technique introduced by Shamir (2017). The technique in Shamir (2017) only proves the regret holding in expectation. In this paper, we extend it to achieve the regret holding with a high probability.
Lemma 8
Letting $\bm{g}_{t}$ be defined Eq. (3) and given $0<\delta<1$, then with a probability at least $1-\delta$, it holds that
$\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}\leq K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right),$ (11)
where $K$ is an absolute constant.
Proof We have
[…]
$\leq$ $\frac{d^{2}}{\alpha^{2}\mu}\cdot K\cdot\frac{(G\alpha)^{2}}{d}\cdot\left(\log T+\log\frac{1}{\delta}\right)$
$=$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right),$
where the last inequality is because of Lemma 19.
Lemma 9
Suppose the loss function $\ell_{t}(\cdot)$ is $\mu$-strongly convex and $G$-Lipschitz continuous. Let the sequence $\{\bm{x}_{t}\}$ be generated by Algorithm 1. Given $0<\delta<1$, then with a probability at least $1-\delta$, it holds that for any $\bm{x}\in(1-\xi)\mathcal{K}$:
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}.$ (12)
Proof Combining Eq. (10) and Eq. (11), we can obtain that
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq$ $\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}$
$\leq$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}.$
Based on above results, we can finally prove Theorem 4.
Proof [Proof of Theorem 4] Combining Lemma 3 and Lemma 9, we can obtain that
$\sum_{t=1}^{T}\frac{\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})+\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})}{2}-\sum_{t=1}^{T}\ell_{t}(\bm{x})$
[…]
$=$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+\frac{(32c_{0}+3)dG^{2}\log\frac{1}{\delta}}{\mu}+\frac{7(d+1)GD\log\frac{1}{\delta}}{3},$
where the second inequality is also because $(1-\xi)\bm{x}\in(1-\xi)\mathcal{K}$ if $\bm{x}\in\mathcal{K}$ and the last inequality is because of $\alpha=\frac{dG\log\frac{1}{\delta}}{T\mu}$ and $\xi=(d+1)\log\frac{1}{\delta}/T$.
4 Conclusion
In this paper, we have addressed a fundamental gap in the theory of bandit online convex optimization by providing the first tight high-probability regret bound for strongly convex functions with two-point feedback. By resolving the open problem posed by Agarwal et al. (2010), we have shown that it is possible to achieve a regret bound of $O(d(\log T+\log(1/\delta))/\mu)$, which is optimal in terms of all key parameters.
Our technical contribution lies in a departure from the standard reduction-based analysis. By developing a direct analytical framework and generalizing the multi-point feedback probabilistic techniques to the high-confidence regime, we have successfully reduced the dimensional dependency from $O(d^{2})$ to $O(d)$. This result bridges the gap between expectation-based bounds and high-probability guarantees in online convex optimization settings.
Appendix A Useful Lemmas
Lemma 10 (Corollary 2.6 of Ledoux (2001))
Letting $\bm{u}\sim\mathbb{S}^{d-1}$, that is, $\bm{u}$ uniformly drawn from the sphere of unit ball in $d$-dimension space, then it holds that
$\Pr\left(|h(\bm{u})-\mathbb{E}[h(\bm{u})]|\geq\tau\right)\leq 2\exp\left(-\frac{c_{1}d\tau^{2}}{L^{2}}\right),$ (13)
where $c_{1}$ is an absolute constant and $L$ is the Lipschitz constant of function $h(\cdot)$.
Lemma 11 (Freedman’s inequality)
Let $\{Y_{t}:t\geq 0\}$ be a real-valued martingale adapted to a filtration $\{\mathcal{F}_{t}\}$, with $Y_{0}=0$. Let $Z_{t}=Y_{t}-Y_{t-1}$ denote the martingale difference sequence. Assume the sequence is uniformly bounded from above:
$Z_{t}\leq R\quad\text{a.s. for all }t\geq 1.$
Define the predictable quadratic variation process $W_{k}$ as:
$W_{t}=\sum_{j=1}^{t}\mathbb{E}[Z_{j}^{2}\mid\mathcal{F}_{j-1}].$
Then, for any $\tau\geq 0$ and $\sigma^{2}>0$:
$\mathbb{P}\left(\exists t:Y_{t}\geq\tau\text{ and }W_{t}\leq\sigma^{2}\right)\leq\exp\left(-\frac{\tau^{2}/2}{\sigma^{2}+R\tau/3}\right).$
Next, we define the Orlicz norm. Then we will introduce definitions and properties of subgaussian and subexponential distribution.
Definition 12
Let $X$ be a random variable, we can define its Orlicz norms $\|X\|_{\psi_{2}}$ and $\|X\|_{\psi_{1}}$ as follows:
$\|X\|_{\psi_{2}}=\inf\left\{\tau>0:\mathbb{E}\exp\left(\frac{X^{2}}{\tau^{2}}\right)\leq 2\right\},$
and
$\|X\|_{\psi_{1}}=\inf\left\{\tau>0:\mathbb{E}\exp\left(\frac{|X|}{\tau}\right)\leq 2\right\}.$ (14)
Lemma 13 (Proposition 2.6.1 of Vershynin (2026))
Let $X$ be a random subgaussian variable. The following properties are equivalent, with the parameters $K_{i}>0$ differing by at most an absolute constant factor.
  1. (i)
    (Tails) There exists $K_{1}>0$ such that
    $\Pr\{|X|\geq\tau\}\leq 2\exp(-\tau^{2}/K_{1}^{2})\quad\text{for all }\tau\geq 0.$ (15)
  2. (ii)
    (Moments) There exists $K_{2}>0$ such that
    $\|X\|_{L^{p}}=(\mathbb{E}|X|^{p})^{1/p}\leq K_{2}\sqrt{p}\quad\text{for all }p\geq 1.$
  3. (iii)
    (MGF of $X^{2}$) There exists $K_{3}>0$ such that
    $\mathbb{E}\exp(X^{2}/K_{3}^{2})\leq 2.$
Moreover, if $\mathbb{E}X=0$ then properties (i)(i)$(iii)$ are equivalent to the following one:
  1. (iv)
    (MGF) There exists $K_{4}>0$ such that
    $\mathbb{E}\exp(\lambda X)\leq\exp(K_{4}^{2}\lambda^{2})\quad\text{for all }\lambda\in\mathbb{R}.$
Above properties can also be represented by Orlicz norm as follows.
Lemma 14 (Proposition 2.6.6 of Vershynin (2026))
Every subgaussian random variable $X$ satisfies the following bounds.
  1. (i)
    (Tails) $\Pr\{|X|\geq\tau\}\leq 2\exp\left(-c\tau^{2}/\|X\|_{\psi_{2}}^{2}\right)$ for all $\tau\geq 0$.
  2. (ii)
    (Moments) $\|X\|_{L^{p}}\leq C\|X\|_{\psi_{2}}\sqrt{p}$ for all $p\geq 1$.
  3. (iii)
    (MGF of $X^{2}$) $\mathbb{E}\exp\left(X^{2}/\|X\|_{\psi_{2}}^{2}\right)\leq 2$.
  4. (iv)
    (MGF) If $\mathbb{E}X=0$ then $\mathbb{E}\exp(\lambda X)\leq\exp\left(C\lambda^{2}\|X\|_{\psi_{2}}^{2}\right)$ for all $\lambda\in\mathbb{R}$.
Here $C,c>0$ are absolute constants. Moreover, up to absolute constant factors, $\|X\|_{\psi_{2}}$ is the smallest possible number that makes each of these statements valid.
Lemma 15 (Proposition 2.8.1 of Vershynin (2026))
Let $X$ be a random subexponential variable. The following properties are equivalent, with the parameters $K_{i}>0$ differing by at most an absolute constant factor.
  1. (i)
    (Tails) There exists $K_{1}>0$ such that
    $\mathbb{P}\{|X|\geq t\}\leq 2\exp(-t/K_{1})\quad\text{for all }t\geq 0.$
  2. (ii)
    (Moments) There exists $K_{2}>0$ such that
    $\|X\|_{L^{p}}=(\mathbb{E}|X|^{p})^{1/p}\leq K_{2}p\quad\text{for all }p\geq 1.$
  3. (iii)
    (MGF of $|X|$) There exists $K_{3}>0$ such that
    $\mathbb{E}\exp(|X|/K_{3})\leq 2.$
Moreover, if $\mathbb{E}X=0$ then properties (i)(i)$(iii)$ are equivalent to the following one:
  1. (iv)
    (MGF) There exists $K_{4}>0$ such that
    $\mathbb{E}\exp(\lambda X)\leq\exp(K_{4}^{2}\lambda^{2})\quad\text{for all }\lambda\text{ such that }|\lambda|\leq\frac{1}{K_{4}}.$
Lemma 16 (Lemma 2.8.5 of Vershynin (2026))
The random variable $X$ is subgaussian if and only if $X^{2}$ is subexponential, and
$\left\|X^{2}\right\|_{\psi_{1}}=\left\|X\right\|^{2}_{\psi_{2}}.$
Appendix B Proofs Related to $\left\|g_{t}\right\|^{2}$
Lemma 17
Letting $\bm{u}\sim\mathbb{S}^{d-1}$ and function $h(\cdot)$ be $L$-Lipschitz, then random variable $X:=h(\bm{u})-\mathbb{E}[h(\bm{u})]$ is a subgaussian random variable which satisfies that
$\left\|X\right\|_{\psi_{2}}\leq C_{1}\frac{L}{\sqrt{d}}.$
Furthermore, the random variable $Y:=X^{2}$ is a sub-exponential random variable with
$\left\|Y\right\|_{\psi_{1}}=\left\|X\right\|^{2}_{\psi_{2}}=\frac{C_{1}^{2}L^{2}}{d}.$ (16)
Letting centered random variable $V=Y-\mathbb{E}[Y]$, then it holds that
$\left\|V\right\|_{\psi_{1}}\leq C_{2}\|Y\|_{\psi_{1}}.$ (17)
Here $C_{1},C_{2}>0$ are absolute constants.
Proof First, by Lemma 10 and Eq. (15), we can obtain that $X$ is a subgaussian random variable. Furthermore, by Eq. (13), item (i) and item (iii) in Lemma 14, we can obtain that
$\left\|X\right\|_{\psi_{2}}\leq C_{1}\frac{L}{\sqrt{d}}.$
By Lemma 16, we can obtain the result about random variable $Y$.
Finally, we have
$\left\|V\right\|_{\psi_{1}}=\left\|Y-\mathbb{E}\left[Y\right]\right\|_{\psi_{1}}\leq C_{2}\|Y\|_{\psi_{1}}.$
Lemma 18
The approximate gradient $\bm{g}_{t}$ defined in Eq. (3) has its norm bounded as
$\|\bm{g}_{t}\|^{2}\leq\frac{d^{2}}{\alpha^{2}}\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2}.$ (18)
Proof Using the definition of $\bm{g}_{t}$, we have
[…]
$\leq\frac{d^{2}}{2\alpha^{2}}\left[\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2}\right$.
$\quad+\left.\left(\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]-\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})\right)^{2}\right]$
$=\frac{d^{2}}{\alpha^{2}}\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2},$
where the last equality is because of the symmetry of the spherical distribution.
Lemma 19
Define random variable $X_{t}=\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}\left[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})\right]$. Given $0<\delta<1$, then with a probability at least $1-\delta$, then it holds that
$\sum_{k=1}^{T}\frac{X_{k}^{2}}{k}\leq K\cdot\frac{(G\alpha)^{2}}{d}\cdot\left(\log T+\log\frac{1}{\delta}\right),$
where $K$ is an absolute constant.
Proof First, we introduce two random variables
$Y_{t}=X_{t}^{2},\qquad\mbox{and}\qquad V_{t}=Y_{t}-\mathbb{E}[Y_{t}].$
Define the natural filtration as the $\sigma$-algebra generated by the random variables $\bm{u}_{t}$’s observed to step $t$:
$\mathcal{F}_{t}:=\sigma(\bm{x}_{1},\bm{u}_{1},\dots,\bm{u}_{t})\mbox{ with }\mathcal{F}_{0}=\sigma(\bm{x}_{1}).$
Defining
$S_{T}=\sum_{k=1}^{T}\frac{V_{k}}{k},$
then, by the definition of $\left\|\cdot\right\|_{\psi_{1}}$ in Eq. (14), items (iii) and item (iv) of Lemma 15, if $\lambda$ satisfies that
$\lambda\leq\frac{c_{1}}{\max_{k}\left\|V_{k}\right\|_{\psi_{1}}/k},$
where $c_{1}$ is an absolute constant, we have
$\mathbb{E}\left[\exp\left(\lambda S_{T}\right)\right]=$ $\mathbb{E}\left[\exp\left(\lambda\sum_{k=1}^{T}\frac{V_{k}}{k}\right)\right]$ (19)
$=$ $\mathbb{E}\left[\mathbb{E}\left[\exp\left(\lambda\sum_{k=1}^{T}\frac{V_{k}}{k}\right)\mid\mathcal{F}_{T-1}\right]\right]$
$=$ $\mathbb{E}\left[\exp\left(\lambda\sum_{k=1}^{T-1}\frac{V_{k}}{k}\right)\mathbb{E}\left[\exp\left(\lambda\frac{V_{T}}{T}\right)\mid\mathcal{F}_{T-1}\right]\right]$
$\leq$ $\mathbb{E}\left[\exp\left(\lambda\sum_{k=1}^{T-1}\frac{V_{k}}{k}\right)\right]\cdot\exp\left(K_{1}\lambda^{2}\frac{\left\|V_{T}\right\|_{\psi_{1}}^{2}}{T^{2}}\right)$
$=$ $\mathbb{E}\left[\exp\left(\lambda S_{T-1}\right)\right]\cdot\exp\left(K_{1}\lambda^{2}\frac{\left\|V_{T}\right\|_{\psi_{1}}^{2}}{T^{2}}\right),$
where $K_{1}$ is an absolute constant and the last inequality is because of item (iv) of Lemma 15.
Using above equation recursively, we can obtain that
$\mathbb{E}\left[\exp\left(\lambda S_{T}\right)\right]\leq\exp\left(\lambda^{2}K_{1}\sum_{k=1}^{T}\frac{\left\|V_{k}\right\|_{\psi_{1}}^{2}}{k^{2}}\right).$
Combining with the Markov’s inequality, we can obtain that
$\Pr\left(\sum_{k=1}^{T}\frac{1}{k}V_{k}\geq\tau\right)\leq\exp(-\lambda\tau)\exp\left(\lambda^{2}K_{1}\sum_{k=1}^{T}\frac{\left\|V_{k}\right\|_{\psi_{1}}^{2}}{k^{2}}\right).$
By choosing
$\lambda=\min\left(\frac{\tau}{2K_{1}\sum_{k=1}^{T}\frac{1}{k^{2}}\|V_{k}\|_{\psi_{1}}^{2}},\frac{c_{1}}{\max_{k=1,\dots,T}\frac{1}{k}\|V_{i}\|_{\psi_{1}}}\right).$
Therefore, we can obtain that
$\Pr\left(\sum_{k=1}^{T}\frac{1}{k}V_{k}\geq\tau\right)\leq\exp\left(-\min\left(\frac{\tau^{2}}{4K_{1}\sum_{k=1}^{T}\frac{1}{k^{2}}\|V_{k}\|_{\psi_{1}}^{2}},\frac{c_{1}\tau}{2\max_{k=1,\dots,T}\frac{1}{k}\|V_{i}\|_{\psi_{1}}}\right)\right).$
Thus, by setting
$\tau=\max\left(\sqrt{4K_{1}\left(\sum_{k=1}^{T}\frac{1}{k^{2}}\|V_{k}\|_{\psi_{1}}^{2}\right)\ln(1/\delta)},\quad\frac{2\max_{k}\frac{1}{k}\|V_{k}\|_{\psi_{1}}}{c_{1}}\ln(1/\delta)\right),$
We can obtain that with a probability at least $1-\delta$, it holds that
$\sum_{k=1}^{T}\frac{1}{k}V_{k}\leq\max\left(\sqrt{4K_{1}\left(\sum_{k=1}^{T}\frac{1}{k^{2}}\|V_{k}\|_{\psi_{1}}^{2}\right)\ln(1/\delta)},\quad\frac{2\max_{k}\frac{1}{k}\|V_{k}\|_{\psi_{1}}}{c_{1}}\ln(1/\delta)\right).$ (20)
Furthermore, it holds that
[…]
where the last inequality is because $\ell_{t}(\bm{x}+\alpha\bm{u})$ is $(G\alpha)$-Lipschitz continuous with respect to $\bm{u}$.
Since $\sum_{k=1}^{T}\frac{1}{k^{2}}<\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$, we can obtain that
$\sum_{k=1}^{T}\frac{1}{k^{2}}\|V_{k}\|_{\psi_{1}}^{2}\leq\frac{\pi^{2}}{6}\cdot\left(\frac{C_{1}^{2}C_{2}(G\alpha)^{2}}{d}\right)^{2}=K_{2}\left(\frac{(G\alpha)^{2}}{d}\right)^{2},$ (21)
where $K_{2}=\frac{\pi^{2}(C_{1}^{2}C_{2})^{2}}{6}$ is an absolute constant.
We also have
$\max_{k=1,\dots,T}\frac{1}{k}\|V_{k}\|_{\psi_{1}}\leq\frac{C_{1}^{2}C_{2}(G\alpha)^{2}}{d}\leq K_{3}\frac{(G\alpha)^{2}}{d},$ (22)
where $K_{3}$ is an absolute constant.
Substitute Eq. (21) and Eq. (22) into Eq. (20) to obtain
$\sum_{k=1}^{T}\frac{1}{k}V_{k}\leq\max\left(\sqrt{4K_{1}K_{2}\left(\frac{(G\alpha)^{2}}{d}\right)^{2}\ln(1/\delta)},\quad\frac{2K_{3}}{c_{1}}\frac{(G\alpha)^{2}}{d}\ln(1/\delta)\right)\leq K_{4}\frac{(G\alpha)^{2}}{d}\ln(1/\delta),$ (23)
where $K_{4}$ is an absolute constant.
Furthermore,
$\sum_{k=1}^{T}\frac{1}{k}\mathbb{E}[Y_{k}]=\sum_{k=1}^{T}\frac{1}{k}\mathbb{E}[X_{k}^{2}]=\sum_{k=1}^{T}\frac{1}{k}\left\|X_{k}\right\|_{2}^{2}$ (24)
[…]
$\leq$ $\frac{K_{5}(G\alpha)^{2}\log T}{d},$
where the first inequality is because of item (ii) of Lemma 14, and $K_{5}$ is an absolute constant. Thus,
$\sum_{k=1}^{T}\frac{1}{k}X_{k}^{2}=$ $\sum_{k=1}^{T}\frac{1}{k}Y_{k}=\sum_{k=1}^{T}\frac{1}{k}V_{k}+\sum_{k=1}^{T}\frac{1}{k}\mathbb{E}[Y_{k}]$
[…]
$\leq$ $K\cdot\frac{(G\alpha)^{2}}{d}\cdot\left(\log T+\log\frac{1}{\delta}\right).$